OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 6x - y = 1 4x - 3y = -11

OpenStudy (anonymous):

a. 4(-6x + 1) - 3y = -11 b. 4(6x + 1) - 3y = -11 c. 4x - 3(6x - 1) = -11 d. 4x - 3(-6x - 1) = -11

OpenStudy (anonymous):

The simplest substitution would be to solve the first equation for \(y\): \[6x-y=1~~\iff~~y=6x-1\] So plugging into the second equation you would have \[\begin{align*} 4x-3\color{red}{y}&=-11\\ 4x-3\color{red}{(6x-1)}&=-11\\ \end{align*}\]

OpenStudy (anonymous):

Thanks